OpenStudy (anonymous):

find 20 airthmetic means between 4 and 67

OpenStudy (vishweshshrimali5):

Lets assume these arithmetic means to be \(\large{a_1, a_2, a_3,..., a_{20}}\). Then, \(\large{4, a_1, a_2, ..., a_{20}, 67}\) must form an AP. Let, the common difference of this AP = d, then, \(\large{a_1 = 4+d}\) Similarly, 67 is 22nd term of this AP. Thus, \[\large{67 = 4 + (22-1)d}\] \[\large{\implies 63 = 21d}\] \[\large{\implies d = \cfrac{63}{21}}\] \[\large{\implies d = 3}\] Thus, \(\large{a_1 = 4 + d = 4+3 = 7}\) \(\large{a_2 = 4 + (3-1)d = 4 + 2* d = 4+2*3 = 10}\) and so on. Thus, the required arithmetic means = 7, 10, 13, 16, 19, ...